A thin, uniform, metal bar, 3 m long and weighing 80 N , is hanging vertically from the ceiling by a frictionless pivot. Suddenly it is struck 1.3 m below the ceiling by a small 4-kg ball, initially traveling horizontally at 14 m/s . The ball rebounds in the opposite direction with a speed of 5 m/s .

Find the angular speed of the bar just after the collision.
Answer #1

bar mass m = 80N / 9.8m/s² = 8.16 kg

Conserve angular momentum:

initial ball mvr = final bar I*ω + final bar mvr

4kg * 14m/s * 1.3m = (8.16kg * (3m)² / 3) * ω + 4kg * -5m/s * 1.3m

98.8 kg·m²/s = 24.48kg·m² * ω

ω = 4.04 rad/s